OpenStudy (anonymous):

Please help! theta= pi/3 part 1: use the equations x=rcos(theta) and y=rsin(theta) to express x and y in terms of r. Two equations after this step. part 2: Use the ratio (y/x) to find the slope of the line defined by this equation.

OpenStudy (ujjwal):

put the values.. That's it!

OpenStudy (anonymous):

Thats what i was thinking it would be for part 1, so it would be x=rcos(pi/3) and y=rsin(pi/3), correct? I don't understand part 2, though.

OpenStudy (anonymous):

http://www.mathsisfun.com/geometry/images/circle-unit-304560.gif use this to find sin and cos of pi/3. and yes that is correct

OpenStudy (ujjwal):

For the second part do as suggested in question.. Divide y by x and substitute the values obtained in part one

OpenStudy (anonymous):

is y=rsin(pi/3) = sqrt3/2 and x=rcos(pi/3)=1/2?

OpenStudy (ujjwal):

you forgot r

OpenStudy (anonymous):

how do I find r?

OpenStudy (ujjwal):

you don't need to find..r will be cancelled out when you divide y by x

OpenStudy (anonymous):

okay so if I did this correctly by dividing the two of them my answer would be sqrt3?

OpenStudy (ujjwal):

yep

OpenStudy (anonymous):

Thanks soooo much! :)

OpenStudy (anonymous):

@ujjwal oh wait, could you help me with one more thing please? :) How would I put this in slope-intercept form using the fact that it passes through the origin?

OpenStudy (ujjwal):

equation of a line passing through origin with slope m is given by y=mx where m is slope second part gives you the value of slope, i.e. m substitute value of m in y=mx and that will be your answer

OpenStudy (anonymous):

thanks again!